OpenStudy (anonymous):

Rewrite using positive exponents (2 / 2 + x^-1) - 3

OpenStudy (mathstudent55):

\(a^{-n} = \dfrac{1}{a^n} \) for \(a \ne 0\).

OpenStudy (anonymous):

thank you :) would you help me apply that formula to the problem please

OpenStudy (mathstudent55):

I'm not sure how to read the problem. Is it \(\dfrac{2}{2 + x^{-1}} - 3\) ?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Then we replace \(x^{-1} \) with what you get using the formula above: \( \dfrac{2}{2 + x^{-1}} - 3\) \( =\dfrac{2}{2 + \frac{1}{x}} - 3 \) Now multiply the numerator and denominator of the fraction by x to get rid of the fraction 1/x in the denominator.

OpenStudy (anonymous):

ok thank you, so the answer would be (2x / 2x + 1) - 3?

OpenStudy (mathstudent55):

\( =\dfrac{2x}{2x+ 1}- 3 \) That is correct. To go further, you need to do the subtraction using a common denominator.

OpenStudy (mathstudent55):

\( =\dfrac{2x}{2x+ 1}- \dfrac{3}{1} \dfrac{2x+1}{2x+1}\) \( =\dfrac{2x}{2x+ 1}- \dfrac{6x+3}{2x+1}\) \( =\dfrac{2x - (6x+3)}{2x+ 1} \) \( =\dfrac{2x - 6x-3}{2x+ 1} \) \( =\dfrac{-4x-3}{2x+ 1} \)

OpenStudy (anonymous):

thank you so much for explaining it so clearly! would you mind helping with one more I have pages of them to do... a^-1 - b^-1 / b^-2 - a^ -2

OpenStudy (mathstudent55):

You mean \( \dfrac{a^{-1} - b^{-1} } {b^{-2} - a^{-2} } \) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would the answer be b^2 - a^2 / a^1 - b^1

OpenStudy (mathstudent55):

I don't know. I haven't solved it yet. Start by changing each negative exponent to a fraction using the formula above.

OpenStudy (mathstudent55):

\( \dfrac{a^{-1} - b^{-1} } {b^{-2} - a^{-2} }\) \( = \dfrac{ \frac{1}{a} - \frac{1}{b} } {\frac{1}{b^2} - \frac{1}{a^2} } \) Now multiply the numerator and denominator by \(a^2b^2\) to get rid of the fractions in the numerator and denominator.

OpenStudy (anonymous):

hmm ok i got ab^2 - a^2b / a^2 - b^2 ...is that right?

OpenStudy (mathstudent55):

\( = \dfrac{ab^2 - a^2b}{a^2 - b^2} \) Yes, I got the same, but we're not done yet. Now factor the numerator and denominator.

OpenStudy (mathstudent55):

\( = \dfrac{ab(b - a)}{(a + b)(a - b)} \) It looks like nothing in the numerator and denominator cancels out, but look at the next step: \( = \dfrac{-ab(a - b)}{(a + b)(a - b)} \) \( = \dfrac{-ab\cancel{(a - b)}}{(a + b)\cancel{(a - b)}} \) \( = \dfrac{-ab}{a + b} \)

OpenStudy (mathstudent55):

This is very useful: a - b = -(b - a) It's easy to show it's true: a - b = -(-a + b) = -(b - a)

OpenStudy (anonymous):

I would never have thought of that, thanks

OpenStudy (mathstudent55):

wlcm